The equation of the line that goes through the points (-7, -2) and (5, 3) can be written in general form Ax + By + C = 0. Find A, B, C

Answer by Alan3354(69443)

Here are some examples:
------------
Write an equation of the line that passes through the given point and satisfies the given condition.
10. (-2, 3) ; parallel to y= 4x + 3
slope m = 4
The slope of lines parallel is the same, m= 4
---------
Use y = mx + b and the point to find b:
3 = -2*4 + b
b = 11
--> y = 4x + 11
------------------
12. (-1, -4) ; perpendicular to y= 2x + 5
Slope = 2
The slope of lines perpendicular is the negative reciprocal, = -1/2
m = -1/2
Do it like #10.
-----------------
Write an equation of the line that passes through the given points.
14. (3,4) , (0,3)
Use determinants:
---------
|+x +y +1|
|+3 +4 +1| = 0
|+0 +3 +1|
------
x(4-3) - y(3-0) +(9-0) = 0
x - 3y = -9
---------------
16. (-5,-4) , (0,11)
Without determinants:
Find the slope, m = diffy/diffx
m = (-4-11)/(-5-0) = 3
Use the slope and either point, and do like #10